During March 2019, Annapolis Corporation recorded $40,000 of costs related to factory overhead.  Alpha's overhead application rate is based on direct labor hours. The preset formula for overhead application estimated that $40,100 would be incurred, and 5,500 direct labor hours would be worked.  During March, 4,750 hours were actually worked. Use this information to determine the amount of factory overhead that was (over) or under applied. (Round answers to the nearest whole dollar. Enter as a positive number if under applied.   Enter as a negative number if over applied.)
Answer #1

Overhead rate = 40100/5500 = 7.29 per labor hour

Applied overhead = 4750*7.29 = 34628

Actual overhead = 40000

Under applied overhead = 40000-34628 = 5372